OpenStudy (anonymous):

Every month, a girl gets allowance. Assume last year she had no money, and kept it up to now. Then she spends 1/2 of her money on clothes, then 1/3 of the remaining money on games, and then 1/4 of the remaining money on toys. After she bought all of that, she had $7777 left. Assuming she only gets money by allowance, how much money does she earn every month?

OpenStudy (sriram):

1/4*x=7777 x=7777*4=31108 money per month=31108/12=2592.33

OpenStudy (anonymous):

how did you get tht?

OpenStudy (sriram):

assume the money wid her=x after spending a fraction of money the money left= 1-fraction for example after spendin 1/2 money left=1-1/2=1/2 hence money left wud be 1/2 of x

OpenStudy (sriram):

then on spending 1/3 money left =(1-1/3) of money left=2/3 * 1/2x=1/3*x

OpenStudy (sriram):

similarly after 1/4 money left=1/3 * 3/4*x=1/4*x

OpenStudy (anonymous):

thankss alot i understand that better than my teacher was trying to explain it

OpenStudy (sriram):

nice complimnt

OpenStudy (anonymous):

no problem she said you hasd to combine your like tems and divise or something??